• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name xiaohui

    Log entry time 00:00:11 on February18,2009

    Entry number 261730

    keyword=shift summary

    shift summary

    16:00 shift start with BCM calibration
    BCM run 1521/20227
    BCM run 1522/20228

    17:30 start with 2uA and BeO target to check the raster and beam position.
    18:00 start raster check
    18:30 ask for a control access
    18:45 fill N2 to reference cell to 10 psig(on display, offset is -12), prepare for N2 data taking
    19:00 back to beam permit
    19:15 beam is back, 2uA on BeO target to check the raster size.
    19:30 take 15uA on N2, Kalyan takes some test runs for bigbite, bigbite wire chamber HV is back on.
    19:40 start N2 reference cell data taking

    run 1529/20236 N2

    20:20 move to 3He and start production (half wave plate is out)

    run 1530/20237 3He

    21:00 vent and vacuumn the reference cell, recorded the offsets for both gauges.

    21:13 ask for control access for some DAQ work.

    23:00 back to beam permit, move target to BeO, check raster size.

    23:20 move target to 3He, start production.

    run 1531/20239 3He


    continue to next shift.